the sum of four consecutive even integers is two more than 5 times the first integer. Find the smallest integer.
Answer by Edwin McCravy(20056)

the sum of four consecutive even integers is 2 more than
5 times the first integer. Find the smallest integer.


first (smallest) even integer = n
second even integer = n+2
third even integer = n+4
fourth even integer = n+6

the sum of four consecutive even integers is 2 more than
5 times the first integer.
n+(n+2)+(n+4)+(n+6) = 5n+2

Find the smallest integer.

Solve for n.

Edwin